# Multi-Momentum Observer Contact Estimation for Bipedal Robots

### v1

@@ Abstract
As bipedal robots become more and more popular in commercial and industrial settings, the ability to control them with a high degree of reliability is critical. To that end, this paper considers how to accurately estimate which feet are currently in contact with the ground so as to avoid improper control actions that could jeopardize the stability of the robot. Additionally, modern algorithms for estimating the position and orientation of a robot’s base frame rely heavily on such contact mode estimates. Dedicated contact sensors on the feet can be used to estimate this contact mode, but these sensors are prone to noise, time delays, damage/yielding from repeated impacts with the ground, and are not available on every robot. To overcome these limitations, we propose a momentum observer based method for contact mode estimation that does not rely on such contact sensors. Often, momentum observers assume that the robot’s base frame can be treated as an inertial frame. However, since many humanoids’ legs represent a significant portion of the overall mass, the proposed method instead utilizes multiple simultaneous dynamic models. Each of these models assumes a different contact condition. A given contact assumption is then used to constrain the full dynamics in order to avoid assuming that either the body is an inertial frame or that a fully accurate estimate of body velocity is known. The (dis)agreement between each model’s estimates and measurements is used to determine which contact mode is most likely using a Markov-style fusion method. The proposed method produces contact detection accuracy of up to 98.44% with a low noise simulation and 77.12% when utilizing data collect on the Sarcos© Guardian® XO® robot (a hybrid humanoid/exoskeleton).
@@ I Introduction
Traditional methods for detecting the contact mode on legged robots make certain assumptions that often do not hold for bipedal (humanoid) robots. These methods typically take one of two approaches. First, a contact sensor, such as a switch or a force-torque sensor, is installed at the robot’s foot so that contact detection becomes a problem of filtering or thresholding a single sensor’s data [1, 2, 3, 4, 5]. This method requires hardware changes and is not ideal for robots requiring robust sensing over long lifespans, as contact sensors are easily damaged due to the impulsive nature of repeated impact with the ground during locomotion. The other method that is often used for contact detection is a thresholding method on estimated contact forces from joint torques [6, 7]. The issue with this method is that it assumes that the robot’s body frame can be treated as an inertial frame and that the legs have significantly lower mass than the body. This assumption does not hold in the case of many bipedal robots as the legs often represent a substantial portion of the robot’s mass, such as the Sarcos© Guardian® XO® shown in Fig. 1 which has close to 50% of total body mass in each leg, and therefore movement of the legs has an appreciable effect on the state of the body.
@@
Figure 1: Left: The Sarcos© Guardian® XO®, a hybrid humanoid robot/exoskeleton on which we demonstrate these algorithms. Right: The estimated external torques on the system during a gait cycle using each separate momentum observer.
@@
This work proposes a new contact estimation system that addresses these challenges through the use of multiple momentum observers that adopt different contact assumptions. We avoid the assumption that the robot’s body frame can be taken as an inertial frame by instead using the world frame as the inertial frame with an assumed stationary foot contact for the stance foot to reduce the order of the dynamics. This assumption allows the 6 constraints at a given foot (in 3D) to remove the need to estimate the base link position, orientation, and velocities. Then, a momentum observer [8, 9, 10, 11, 12] is applied to the constrained dynamics of the system in each contact mode. By comparing the results of momentum observers on each of these assumptions, as well as monitoring the relative velocity between the feet, we obtain a probability of each contact condition (including dual support).
@@
The proposed approach is thus a novel combination of reduced order dynamics, momentum observers, and Markov models which enables contact detection in large scale bipedal robots without (i) direct contact sensors, (ii) using the second derivative of any measurement, (iii) precise knowledge of the body velocity, or (iv) assuming that the legs are much lighter than the body or moving much faster than the body. The knowledge of contact states provided by this algorithm will enable the use of a variety of full body state estimators that rely on contact knowledge.
@@ III-A Momentum observers
where $\{q,\dot{q},\ddot{q}\}$ represent the generalized coordinates of the system (i.e., joint angles + base position/orientation) and their derivatives, $M$ is the generalized momentum matrix, $C$ is the Coriolis matrix, and $G$ is the gravity vector. Modeled constraint forces ($\lambda$) act on the system through the constraint matrix $A$, while applied motor torques ($\tau_{\text{mot}}$) and unmodeled/external torques ($\tau_{\text{ext}}$) act through the selector matrix ($B^{T}\tau\equiv[0_{6\times 1}^{T}~{}\tau^{T}]^{T}$) that simply ensures the floating base degrees-of-freedom (DOFs) are not actuated. Trivially, it is possible to calculate the external torques with measurements/estimates of $q$ (and its derivatives), $\lambda$, and $\tau_{\text{mot}}$ utilizing the inverse dynamics approach. But rather than relying on noisy acceleration data, we instead consider the derivative of the system’s generalized momentum $p=M\dot{q}$:
@@ III-B Constrained Dynamics
$&H^{T}MH\ddot{y}+\left(H^{T}CH+H^{T}M\dot{H}\right)\dot{y}\\ &\hskip 20.00003pt+H^{T}G+H^{T}A^{T}\lambda=H^{T}B^{T}(\tau_{\text{mot}}+\tau_{\text{ext}})\end{aligned}$ (7a)
@@ III-E Model Fusion
$\pi_{i,\text{Dual}}=\text{sgm}\left(\frac{||\hat{\tau}_{i,\text{ext}}||_{2}-\tau_{t}}{\tau_{b}}\right)$ (13)
@@
$\pi_{\text{Dual},i}=\text{sgm}\left(\frac{v_{\text{rel}}-v_{t}}{v_{b}}\right)\frac{||\hat{\tau}_{j,\text{ext}}||_{2}}{||\hat{\tau}_{i,\text{ext}}||_{2}+||\hat{\tau}_{j,\text{ext}}||_{2}}$ (14)
where, for each $i\in\{l,r\}$, $j\in\{l,r\}\backslash\{i\}$ and $\pi_{Dual,i}$ is the transition probability from dual support to the given single support mode.
@@
To select the active contact mode estimate, we take the most likely mode at a given time. This may sometimes result in selecting an active mode that is estimated to have less than a 50% likelihood. In practice, there are many ways to handle this information. The likelihoods themselves could be fed into other estimation methods as a means of representing how much certainty should be associated with dynamic constraints from a contact mode. Alternatively, an additional mode that represents an uncertain contact condition could be added in order to prevent switching when uncertainty is more likely. For the work proposed here, the contact mode estimate is only allowed to transition to a new contact mode once a threshold of belief in a given mode is crossed (e.g., $>50$% likelihood for a given mode is required to switch to that mode).
@@ IV-A Simulated Walker
Figure 2: The five link walker model we use in simulation to evaluate the algorithm. The floating base location is defined as a position and angle relative to the vertical. The link angles are defined relative to straight hip and straight knee positions.
@@ IV-B Exoskeleton Walker
For this portion of the work, we utilize data collected from a Sarcos© Guardian® XO® robot (a hybrid humanoid/exoskeleton), which is shown in Fig. 1. This robot has a mass of 150kg and much of that mass is distributed in the limbs, so the assumption that the body link is an inertial frame cannot be adopted. This data was collected during a trial in which the XO walked as a humanoid on its own without a person inside controlling it. The measurements used for this experiment are the measured joint torques (as opposed to the commanded), the orientation of the body via an IMU, and the joint positions via encoders.
All of the mode estimates on the XO are compared against hand-labeled contact modes derived from a video of the corresponding trial. A confusion matrix comparing the estimated mode to the mode extracted from the video is shown in Fig. 6. The associated estimates from the trial are shown as a time series in Fig. 7. The corresponding mode accuracy for this system is 77.12%, compared to 70.15% from the planned contact sequence. Directly utilizing the most likely contact mode for a given time without the probabilistic transition model results in a similar accuracy of 76.10%, however it predicts twice as many transitions than truly occur compared to 27.5% fewer than true transitions predicted by the transition model. While this is not a perfect model of the transitions, it is preferable to have a stable estimate of the current contact state for control purposes. Some of the misalignment can be explained by the difference in the estimation update rate and the camera frame rate. Additionally, due to the hand-labeled nature of the ground truth, there are orientations in which it is difficult to discern the true contact conditions. However, the general timing and gait patterns align well between the estimate and the values derived from the associated video. To improve the practical performance, the cutoff values for the sigmoids can be tuned to either over- or under-estimate dual support depending on the context in which the estimation is being used. Additionally, this should be fused with more information, such as estimated ground heights from a visual or lidar sensor. However, for this work we sought to make this estimation agnostic to the profile of the ground.

### v2

@@ Abstract
As bipedal robots become more and more popular in commercial and industrial settings, the ability to control them with a high degree of reliability is critical. To that end, this paper considers how to accurately estimate which feet are currently in contact with the ground so as to avoid improper control actions that could jeopardize the stability of the robot. Additionally, modern algorithms for estimating the position and orientation of a robot’s base frame rely heavily on such contact mode estimates. Dedicated contact sensors on the feet can be used to estimate this contact mode, but these sensors are prone to noise, time delays, damage/yielding from repeated impacts with the ground, and are not available on every robot. To overcome these limitations, we propose a momentum observer based method for contact mode estimation that does not rely on such contact sensors. Often, momentum observers assume that the robot’s base frame can be treated as an inertial frame. However, since many humanoids’ legs represent a significant portion of the overall mass, the proposed method instead utilizes multiple simultaneous dynamic models. Each of these models assumes a different contact condition. A given contact assumption is then used to constrain the full dynamics in order to avoid assuming that either the body is an inertial frame or that a fully accurate estimate of body velocity is known. The (dis)agreement between each model’s estimates and measurements is used to determine which contact mode is most likely using a Markov-style fusion method. The proposed method produces contact detection accuracy of up to 98.44% with a low noise simulation and 77.12% when utilizing data collect on a hybrid humanoid/exoskeleton.
@@ I Introduction
Traditional methods for detecting the contact mode on legged robots make certain assumptions that often do not hold for bipedal (humanoid) robots. These methods typically take one of two approaches. First, a contact sensor, such as a switch or a force-torque sensor, is installed at the robot’s foot so that contact detection becomes a problem of filtering or thresholding a single sensor’s data [3, 13, 9, 8, 12]. This method requires hardware changes and is not ideal for robots requiring robust sensing over long lifespans, as contact sensors are easily damaged due to the impulsive nature of repeated impact with the ground during locomotion. The other method that is often used for contact detection is a thresholding method on estimated contact forces from joint torques [4, 18]. The issue with this method is that it assumes that the robot’s body frame can be treated as an inertial frame and that the legs have significantly lower mass than the body. This assumption does not hold in the case of many bipedal robots as the legs often represent a substantial portion of the robot’s mass, such as the hybrid humanoid/exoskeleton shown in Fig. 1 which has close to 50% of total body mass in the legs, and therefore movement of the legs has an appreciable effect on the state of the body.
@@
Figure 1: Left: The hybrid humanoid robot/exoskeleton on which we demonstrate these algorithms. Right: The estimated external torques on the system during a gait cycle using each separate momentum observer.
@@
This work proposes a new contact estimation system that addresses these challenges through the use of multiple momentum observers that adopt different contact assumptions. We avoid the assumption that the robot’s body frame can be taken as an inertial frame by instead using the world frame as the inertial frame with an assumed stationary foot contact for the stance foot to reduce the order of the dynamics. The proposed method focuses on bipedal robots as robots with more legs often have legs representing lower fractions of the total mass, enabling the use of the body as an inertial frame, or have more legs in contact at a given time, allowing quasi-static analysis. This assumption allows the 6 constraints at a given foot (in 3D) to remove the need to estimate the base link position, orientation, and velocities. Then, a momentum observer [5, 6, 7, 1, 19] is applied to the constrained dynamics of the system in each contact mode. By comparing the results of momentum observers on each of these assumptions, as well as monitoring the relative velocity between the feet, we obtain a probability of each contact condition (including dual support).
@@
The proposed approach is thus a novel combination of reduced order dynamics, momentum observers, and Markov models which enables contact detection in large scale bipedal robots without (i) direct contact sensors as in [3, 13, 9, 8, 12], (ii) using the second derivative of any measurement or precise knowledge of the body velocity as enabled by [5, 6, 7, 1, 19], or (iii) assuming that the legs are much lighter than the body or moving much faster than the body as in [4, 18]. The knowledge of contact states provided by this algorithm will enable the use of a variety of full body state estimators that rely on contact knowledge.
@@ III-A Momentum observers
where $\{q,\dot{q},\ddot{q}\}$ represent the generalized coordinates of the system (i.e., joint angles + base position/orientation) and their derivatives, $M$ is the mass matrix, $C$ is the Coriolis matrix, and $G$ is the gravity vector. Modeled constraint forces ($\lambda$) act on the system through the constraint matrix $A$, while applied motor torques ($\tau_{\text{mot}}$) and unmodeled/external torques ($\tau_{\text{ext}}$) act through the selector matrix ($B^{T}\tau\equiv[0_{6\times 1}^{T}~\tau^{T}]^{T}$) that simply ensures the floating base degrees-of-freedom (DOFs) are not actuated. Trivially, it is possible to calculate the external torques with measurements/estimates of $q$ (and its derivatives), $\lambda$, and $\tau_{\text{mot}}$ utilizing the inverse dynamics approach. But rather than relying on noisy acceleration data, we instead consider the derivative of the system’s generalized momentum $p=M\dot{q}$:
@@ III-B Constrained Dynamics
$\begin{aligned} &H^{T}MH\ddot{y}+\left(H^{T}CH+H^{T}M\dot{H}\right)\dot{y}\\ &\hskip 20.00003pt+H^{T}G+H^{T}A^{T}\lambda=H^{T}B^{T}(\tau_{\text{mot}}+\tau_{\text{ext}})\end{aligned}$ (7a)
@@ III-E Model Fusion
[…]
@@
[…]
@@
To select the active contact mode estimate, we take the most likely mode at a given time. This may sometimes result in selecting an active mode that is estimated to have less than a 50% likelihood. In practice, there are many ways to handle this information. The likelihoods themselves could be fed into other estimation methods as a means of representing how much certainty should be associated with dynamic constraints from a contact mode. Alternatively, an additional mode that represents an uncertain contact condition could be added in order to prevent switching when uncertainty is more likely. For the work proposed here, the contact mode estimate is only allowed to transition to a new contact mode once a threshold of belief in a given mode is crossed (e.g., >50>50% likelihood for a given mode is required to switch to that mode).
@@ IV-A Simulated Walker
Figure 2: The five link walker model we use in simulation to evaluate the algorithm. The floating base location is defined as a position ($x_{b},y_{b}$) and angle ($\theta_{b}$) relative to the vertical. The link angles are defined relative to straight hip ($\theta_{1},\theta_{3}$) and straight knee ($\theta_{2},\theta_{4}$) positions.
@@ IV-B Exoskeleton Walker
For this portion of the work, we utilize data collected from a hybrid humanoid/exoskeleton, which is shown in Fig. 1. This robot has a mass of 150kg and much of that mass is distributed in the limbs, so the assumption that the body link is an inertial frame cannot be adopted. This data was collected during a trial in which the system walked as a humanoid on its own without a person inside controlling it. The measurements used for this experiment are the measured joint torques (as opposed to the commanded), the orientation of the body via an IMU, and the joint positions via encoders.
All of the mode estimates on the system are compared against hand-labeled contact modes derived from a video of the corresponding trial. A confusion matrix comparing the estimated mode to the mode extracted from the video is shown in Fig. 6. The associated estimates from the trial are shown as a time series in Fig. 7. The corresponding mode accuracy for this system is 77.12%, compared to 70.15% accuracy of the feedforward planned contact sequence from the motion planner. Directly utilizing the most likely contact mode for a given time without the probabilistic transition model results in a similar accuracy of 76.10%, however it predicts twice as many transitions than truly occur compared to 27.5% fewer than true transitions predicted by the transition model. While this is not a perfect model of the transitions, it is preferable to have a stable estimate of the current contact state for control purposes. Some of the misalignment can be explained by the difference in the estimation update rate and the camera frame rate. Additionally, due to the hand-labeled nature of the ground truth, there are orientations in which it is difficult to discern the true contact conditions. However, the general timing and gait patterns align well between the estimate and the values derived from the associated video. To improve the practical performance, the cutoff values for the sigmoids can be tuned to either over- or under-estimate dual support depending on the context in which the estimation is being used. Additionally, to reach higher accuracy values for use in real time control, this should be fused with more information, such as estimated ground heights from a visual or lidar sensor. However, for this work we sought to make this estimation agnostic to the profile of the ground.